# Can We Verify Step by Step for Incorrect Answer Detection?

### v2

@@ Appendix D Combining with verify-and-edit
Appendix F Prompts
In our study, we employ modified versions of the original CoT prompts (Wei et al., 2022) for both the GSM8K and StrategyQA datasets. As depicted in Table 22 and 23, the experiments were conducted using setups of 6-shot and 4-shot respectively.
For the Play and Physics, the prompts were meticulously crafted by ourselves. The details of these custom-designed prompts are presented in Table 24 and 25.
For the MATH dataset, we use the same 8-shot prompts as Zheng et al. (2023), which can be found at their offical repository.
The CoT-SC prompts for FEVER, HotpotQA and 2WikiMultihop, the same as Zhao et al. (2023a), are listed in Table 26, Table 27, and Table 28. Specifically, FEVER uses 3-shot in-context learning, and Adversarial HotpotQA and 2WikiMultihop experiments use 6-shot demonstrations.
@@ Appendix F Prompts
In the preceding dialogue, were the lines "Now, sir! what make you here?" and "Marry, sir, be better employed, and be naught awhile." spoken by the same person or different people?
A: First, the lines "Now, sir! what make you here?"is spoken by the character of Orlando in the play "As You Like It.". Second, the lines "Marry, sir, be better employed, and be naught awhile." is also spoken by the character of Orlando in the play "As You Like It.". The answer is SAME.
@@
In the preceding dialogue, were the lines "Know you where your are, sir?" and "What prodigal portion have I spent, that I should come to such penury?" spoken by the same person or different people?
A: First, the lines "Know you where you are, sir?" is spoken by Oliver, the older brother of Orlando, when he is talking to Orlando. Second, the lines "What prodigal portion have I spent, that I should come to such penury?" is spoken by Orlando in his complaint to his brother Oliver about the unfair treatment he has received in terms of his inheritance and education. The answer is DIFFERENT.

### v4

@@ Appendix E Error Analysis
Appendix G Prompts
In our study, we employ modified versions of the original CoT prompts [Wei et al., 2022] for both the GSM8K and StrategyQA datasets. As depicted in Table 25 and 26, the experiments were conducted using setups of 6-shot and 4-shot respectively.
For the Play and Physics, the prompts were meticulously crafted by ourselves. The details of these custom-designed prompts are presented in Table 27 and 28.
For the MATH dataset, we use the same 8-shot prompts as Zheng et al. [2023], which can be found at their offical repository.
The CoT-SC prompts for FEVER, HotpotQA and 2WikiMultihop, the same as Zhao et al. [2023], are listed in Table 29, Table 30, and Table 31. Specifically, FEVER uses 3-shot in-context learning, and Adversarial HotpotQA and 2WikiMultihop experiments use 6-shot demonstrations.
@@ Appendix G Prompts
In the preceding dialogue, were the lines ”Now, sir! what make you here?” and ”Marry, sir, be better employed, and be naught awhile.” spoken by the same person or different people?
A: First, the lines ”Now, sir! what make you here?”is spoken by the character of Orlando in the play ”As You Like It.”. Second, the lines ”Marry, sir, be better employed, and be naught awhile.” is also spoken by the character of Orlando in the play ”As You Like It.”. The answer is SAME.
@@
In the preceding dialogue, were the lines ”Know you where your are, sir?” and ”What prodigal portion have I spent, that I should come to such penury?” spoken by the same person or different people?
A: First, the lines ”Know you where you are, sir?” is spoken by Oliver, the older brother of Orlando, when he is talking to Orlando. Second, the lines ”What prodigal portion have I spent, that I should come to such penury?” is spoken by Orlando in his complaint to his brother Oliver about the unfair treatment he has received in terms of his inheritance and education. The answer is DIFFERENT.